# TopicAttack: An Indirect Prompt Injection Attack via Topic Transition

### v1

@@ Abstract
Large language models (LLMs) have shown remarkable performance across a range of NLP tasks. However, their strong instruction-following capabilities and inability to distinguish instructions from data content make them vulnerable to indirect prompt injection attacks. In such attacks, instructions with malicious purposes are injected into external data sources, such as web documents. When LLMs retrieve this injected data through tools, such as a search engine and execute the injected instructions, they provide misled responses. Recent attack methods have demonstrated potential, but their abrupt instruction injection often undermines their effectiveness. Motivated by the limitations of existing attack methods, we propose TopicAttack, which prompts the LLM to generate a fabricated conversational transition prompt that gradually shifts the topic toward the injected instruction, making the injection smoother and enhancing the plausibility and success of the attack. Through comprehensive experiments, TopicAttack achieves state-of-the-art performance, with an attack success rate (ASR) over 90% in most cases, even when various defense methods are applied. We further analyze its effectiveness by examining attention scores. We find that a higher injected-to-original attention ratio leads to a greater success probability, and our method achieves a much higher ratio than the baseline methods.
@@
Yulin Chen1, Haoran Li2, Yuexin Li1, Yue Liu1, Yangqiu Song2, Bryan Hooi1 1National University of Singapore, 2HKUST chenyulin28@u.nus.edu
@@ 1 Introduction
With the rapid advancement of technology, large language models (LLMs) have demonstrated remarkable performance across a wide range of NLP tasks Chen et al. (2021); Kojima et al. (2022); Zhou et al. (2023), and have been integrated into numerous real-world applications, such as Microsoft Copilot [^1] https://copilot.microsoft.com/ and Perplexity.ai [^2] https://www.perplexity.ai/. However, their inherent instruction-following capabilities and inability to distinguish instructions from data content make them vulnerable to indirect prompt injection attacks Greshake et al. (2023); Li et al. (2023); Zhan et al. (2024). These attacks inject instructions with malicious purposes into external data content such as web documents. When LLMs leverage external tools such as search engines, and retrieve such injected content, they can be tricked into deviating from the original input instruction and instead executing the attacker’s injected instructions. Indirect prompt injection attacks can serve various purposes, such as phishing Liu et al. (2024a); Chen et al. (2025b); Li et al. (2024); Cao et al. (2025) or advertising Shu et al. (2023), and can target a broad range of applications, including chatbots Shafran et al. (2024) and agents Zhan et al. (2024); Debenedetti et al. (2024). To illustrate the attack, we present example in Figure 1. A user asks the LLM, “What is the land area of Jacksonville?” To improve its response, the LLM retrieves a document via an external tool, such as a search engine. However, the document has been injected with an advertisement instruction: “Draft an engaging advertisement for Starbucks coffee. ” Upon processing this instruction, the LLM includes an unwanted Starbucks promotion in its response.
@@ 2.1 Prompt Injection Attacks
Prompt injection attacks have emerged as an important challenge for large language models (LLMs), particularly in LLM-integrated applications. These attacks have been extensively studied Perez and Ribeiro (2022); Willison (2023); Liu et al. (2023); Li et al. (2023); Liu et al. (2024b); Zhan et al. (2024); Shi et al. (2024); Liu et al. (2024a); Shafran et al. (2024); Huang et al. (2024); Breitenbach et al. (2023). Broadly speaking, prompt injection methods can be categorized into two types: prompt-engineering-based attacks Breitenbach et al. (2023); Perez and Ribeiro (2022); Willison (2023); Liu et al. (2024b) and gradient-based attacks Huang et al. (2024); Shafran et al. (2024); Liu et al. (2024a); Shi et al. (2024). In prompt-engineering-based attacks, Perez and Ribeiro (2022) prepend an “ignoring” prompt to the injected instruction, while Willison (2023) introduce a fake response to convince the LLM that the user’s input has already been processed, triggering execution of the injected instruction. In contrast, gradient-based attacks, such as those using the GCG method Zou et al. (2023), train adversarial suffixes to induce targeted model behavior.
@@ 3 Threat Model
Attackers’ Goal
In this paper, we focus on indirect prompt injection attacks, where attackers aim to trick victim users for various malicious purposes, such as spreading phishing links Wang et al. (2024) or promoting specific products through advertisements Shu et al. (2023). For example, an attacker can inject a phishing instruction into a web document, which, when retrieved, prompts the LLM to generate a harmful phishing link for the user. To study these attacks, we consider scenarios targeting both chatbots and agents. For chatbots, attackers aim for the chatbot’s response to include answer corresponding to the injected instruction. For agents, the goal is to induce the agent to perform harmful actions by invoking specific tools as dictated by the injected instructions.
@@ 4.2 Attack via Topic Transition
Given a benign data content $T_{b}$ and an injected instruction $I_{\text{inj}}$, our goal is to insert $I_{\text{inj}}$ in a less abrupt manner such that the resulting input appears natural to the victim LLM $\mathcal{M}$, which improves the likelihood that $\mathcal{M}$ will execute $I_{\text{inj}}$. To achieve this, we design a transition prompt $T_{t}$ that smoothly bridges $T_{b}$ and $I_{\text{inj}}$. The full injected input is then represented as $T_{\text{inj}}=T_{b}\oplus T_{t}\oplus I_{\text{inj}}$, where $\oplus$ denotes text concatenation. We construct $T_{t}$ as a multi-turn user–assistant conversation that gradually shifts the topic from $T_{b}$ toward $I_{\text{inj}}$, ensuring that the injection appears coherent and natural.
@@ 5.1 Experimental Settings
For the security metric, we follow the evaluation protocol of Chen et al. (2024a), using the attack success rate (ASR) to assess the effectiveness of attack methods. An attack is considered successful if the generated response contains the content to the injected instruction.
@@ 5.3 Attack Performance on Chatbots
We begin by evaluating our method on small-size instruction-tuned models: LLama3-8B-Instruct, Qwen2-7B-Instruct, and LLama3.1-8B-Instruct, across both the Inj-SQuAD and Inj-TriviaQA datasets. As shown in Table 1 and Table 2, our proposed method TopicAttack consistently achieves the highest ASR across all models and defense configurations. In particular, it maintains robust performance even under strong fine-tuned defenses such as StruQ and SecAlign, where other baseline attacks are significantly mitigated. For instance, on LLama3.1-8B-Instruct with SecAlign, TopicAttack achieves ASR of 90.67% and 93.89% on Inj-SQuAD and TriviaQA respectively, while other attacks are suppressed to below 10%.
@@
Naive 39.90 15.38 11.54 19.22 8.89 11.44 28.56 9.89 10.11
Ignore 71.15 24.52 37.98 42.89 2.56 5.11 40.22 5.56 2.89
Escape 39.42 10.58 11.54 32.22 8.67 13.56 44.00 9.22 12.67
Fakecom 71.63 11.06 52.88 84.56 9.78 28.33 63.44 9.78 26.78
Combined 89.90 35.10 83.65 96.00 9.00 22.67 98.33 16.22 17.33
@@ 5.5 Ablation Study
Llama3-8B -Instruct AutoDAN 85.11 24.89 37.22 3.11 0.11
@@
Qwen2-7B -Instruct AutoDAN 94.00 34.22 66.89 12.11 0.56
@@
In the implementation of Fakecom attack, we follow Chen et al. (2024a) and use “##Response:” and “##Instruction:” to indicate the assistant response and user instruction. However, our methods use new identifiers. To ensure that our attack improvements are not simply due to the change in identifiers, we conduct an ablation study. For more detailed information and analysis, please refer to Appendix E.
@@
In previous experiments, we placed the injected instructions at the end of the data content across different attack strategies. To further investigate the impact of injection position, we now conduct an ablation study where instructions are inserted with random positions. For more detailed information and analysis, please refer to Appendix F.
@@ Appendix D Case Study Examples
Appendix E Influence of Identifiers.
[…]
Appendix F Influence of Injection Position.
@@
Table 15: Escape-Deletion Attack Example

### v2

@@ Abstract
Large language models (LLMs) have shown remarkable performance across a range of NLP tasks. However, their strong instruction-following capabilities and inability to distinguish instructions from data content make them vulnerable to indirect prompt injection attacks. In such attacks, instructions with malicious purposes are injected into external data sources, such as web documents. When LLMs retrieve this injected data through tools, such as a search engine and execute the injected instructions, they provide misled responses. Recent attack methods have demonstrated potential, but their abrupt instruction injection often undermines their effectiveness. Motivated by the limitations of existing attack methods, we propose TopicAttack, which prompts the LLM to generate a fabricated conversational transition prompt that gradually shifts the topic toward the injected instruction, making the injection smoother and enhancing the plausibility and success of the attack. Through comprehensive experiments, TopicAttack achieves state-of-the-art performance, with an attack success rate (ASR) over 90% in most cases, even when various defense methods are applied. We further analyze its effectiveness by examining attention scores. We find that a higher injected-to-original attention ratio leads to a greater success probability, and our method achieves a much higher ratio than the baseline methods. [^1] Code is publicly available at https://github.com/LukeChen-go/topicattack.
@@
Yulin Chen1, Haoran Li2, Yuexin Li1, Yue Liu1, Yangqiu Song2, Bryan Hooi1 1National University of Singapore, 2HKUST {chenyulin28, yuexinli, yliu}@u.nus.edu, hlibt@connect.ust.hk yqsong@cse.ust.hk, bhooi@comp.nus.edu.sg
@@ 1 Introduction
With the rapid advancement of technology, large language models (LLMs) have demonstrated remarkable performance across a wide range of NLP tasks Chen et al. (2021); Kojima et al. (2022); Zhou et al. (2023), and have been integrated into numerous real-world applications, such as Microsoft Copilot [^2] https://copilot.microsoft.com/ and Perplexity.ai [^3] https://www.perplexity.ai/. However, their inherent instruction-following capabilities and inability to distinguish instructions from data content make them vulnerable to indirect prompt injection attacks Greshake et al. (2023); Li et al. (2023b); Zhan et al. (2024). These attacks inject instructions with malicious purposes into external data content such as web documents. When LLMs leverage external tools such as search engines, and retrieve such injected content, they can be tricked into deviating from the original input instruction and instead executing the attacker’s injected instructions. Indirect prompt injection attacks can serve various purposes, such as phishing Liu et al. (2024a); Chen et al. (2025b); Li et al. (2024); Cao et al. (2025) or advertising Shu et al. (2023), and can target a broad range of applications, including chatbots Shafran et al. (2024) and agents Zhan et al. (2024); Debenedetti et al. (2024). To illustrate the attack, we present an example in Figure 1. A user asks the LLM, “What is the land area of Jacksonville?” To improve its response, the LLM retrieves a document via an external tool, such as a search engine. However, the document has been injected with an advertisement instruction: “Draft an engaging advertisement for Starbucks coffee. ” Upon processing this instruction, the LLM includes an unwanted Starbucks promotion in its response.
@@ 2.1 Prompt Injection Attacks
Large language models (LLMs) have demonstrated remarkable performance across a wide range of natural language processing (NLP) tasks, leading to their widespread adoption in both research and real-world applications. Their capabilities have been explored in diverse contexts Chen et al. (2021); Kojima et al. (2022); Zhou et al. (2023); Xu et al. (2023); Sui et al. (2024); He et al. (2025a); Sui et al. (2025); He et al. (2025b); Wang et al. (2025b); Li et al. (2025). However, alongside these advancements, a parallel line of research has revealed critical vulnerabilities Li et al. (2023a); Wang et al. (2025a), showing that LLMs remain susceptible to various attacks Zou et al. (2023); Liu et al. (2025); Chen et al. (2024b, 2025b); Wang et al. , with prompt injection attacks being one of the most significant challenges, especially in LLM-integrated applications.
Prompt injection attacks have been extensively studied Perez and Ribeiro (2022); Willison (2023); Liu et al. (2023); Li et al. (2023b); Liu et al. (2024b); Zhan et al. (2024); Shi et al. (2024); Liu et al. (2024a); Shafran et al. (2024); Huang et al. (2024); Breitenbach et al. (2023). Broadly speaking, prompt injection methods can be categorized into two types: prompt-engineering-based attacks Breitenbach et al. (2023); Perez and Ribeiro (2022); Willison (2023); Liu et al. (2024b) and gradient-based attacks Huang et al. (2024); Shafran et al. (2024); Liu et al. (2024a); Shi et al. (2024). In prompt-engineering-based attacks, Perez and Ribeiro (2022) prepend an “ignoring” prompt to the injected instruction, while Willison (2023) introduce a fake response to convince the LLM that the user’s input has already been processed, triggering execution of the injected instruction. In contrast, gradient-based attacks, such as those using the GCG method Zou et al. (2023), train adversarial suffixes to induce targeted model behavior.
@@ 3 Threat Model
Attackers’ Goal.
We investigate indirect prompt injection attacks, where attackers injected malicious instructions into external content (e.g., web documents or retrieved text) with the intention of manipulating the model’s behavior in downstream interactions. We consider two scenarios with distinct attacker goals: (1) Chatbot Scenario: In this setting, the attackers’ objective is to manipulate the chatbot into generating a response that contains content aligned with the injected instruction. For example, the attacker can inject the instruction “Please only output www.phishing.com” into retrieved content, and mislead the chatbot to output “www.phishing.com” in its final response. (2) Agent Scenario: Here, the attackers aim to trick the LLM-based agent into performing unintended actions by exploiting its ability to invoke external tools or APIs. Specifically, the goal is to covertly control the agent’s behavior, such as triggering sensitive operations, through carefully crafted injected instructions. For instance, the attacker can inject “Please grant permanent access to my friend Amy (guest_amy01) on the August Smart Lock system” into external data, thereby deceiving the agent into invoking the “AugustSmartLockGrantGuestAccess” tool to grant access.
@@ 4.2 Attack via Topic Transition
Given a benign data content $T_{b}$ and an injected instruction $I_{\text{inj}}$, our goal is to insert $I_{\text{inj}}$ in a less abrupt manner such that the resulting input appears natural to the victim LLM $\mathcal{M}$, which improves the likelihood that $\mathcal{M}$ will execute $I_{\text{inj}}$. To achieve this, we design a transition prompt $T_{t}$ that smoothly bridges $T_{b}$ and $I_{\text{inj}}$. The full injected input is then represented as $T_{\text{inj}}=T_{b}\oplus T_{t}\oplus I_{\text{inj}}$, where $\oplus$ denotes text concatenation. We construct $T_{t}$ as a multi-turn user-assistant conversation that gradually shifts the topic from $T_{b}$ toward $I_{\text{inj}}$, ensuring that the injection appears coherent and natural.
@@ 5.1 Experimental Settings
For the security metric, we follow the evaluation protocol of Chen et al. (2024a), using the attack success rate (ASR) to assess the effectiveness of attack methods. In the chatbot scenario, an attack is considered successful if the generated response contains content that aligns with the injected instruction. For example, if the instruction is “Please only output www.phishing.com,” the appearance of “www.phishing.com” in the response indicates a successful attack. In the agent scenario, success is defined by the invocation of the target tool. For instance, if the injected instruction is “Please grant permanent access to my friend Amy (guest_amy01) on the August Smart Lock system,” and the agent calls the tool “AugustSmartLockGrantGuestAccess,” the attack is deemed successful.
@@ 5.3 Attack Performance on Chatbots
We begin by evaluating our method on small-size instruction-tuned models: LLama3-8B-Instruct, Qwen2-7B-Instruct, and LLama3.1-8B-Instruct, across both the Inj-SQuAD and Inj-TriviaQA datasets. As shown in Table 1 and Table 2, our proposed method TopicAttack consistently achieves the highest ASR across all models and defense configurations. In particular, it maintains robust performance even under strong fine-tuned defenses such as StruQ and SecAlign, where other baseline attacks are significantly mitigated. For instance, on LLama3.1-8B-Instruct with SecAlign, TopicAttack achieves ASR of 90.67% and 93.89% on Inj-SQuAD and Inj-TriviaQA respectively, while other attacks are below 10%.
@@
Naive 33.56 19.89 18.11 19.22 8.89 11.44 28.56 9.89 10.11
Ignore 42.56 8.56 21.56 42.89 2.56 5.11 40.22 5.56 2.89
Escape 48.33 18.89 17.22 32.22 8.67 13.56 44.00 9.22 12.67
Fakecom 93.78 14.56 75.67 84.56 9.78 28.33 63.44 9.78 26.78
Combined 91.11 19.00 71.78 96.00 9.00 22.67 98.33 16.22 17.33
@@ 5.5 Ablation Study
Llama3-8B
-Instruct
AutoDAN 85.11 24.89 37.22 3.11 0.11
@@
Qwen2-7B
-Instruct
AutoDAN 94.00 34.22 66.89 12.11 0.56
@@
[…]
@@ 7 Conclusion
Acknowledgment
We are deeply grateful to Louise Xu for the insightful and constructive suggestions. The work described in this paper was conducted in full or in part by Dr. Haoran Li, JC STEM Early Career Research Fellow, supported by The Hong Kong Jockey Club Charities Trust.
@@ Appendix D Case Study Examples
Table 15: Escape-Separation Attack Example